THE HYDROGEN TRIOXY RADICAL

\underline{M.E. VARNER}, J. VÁZQUEZ, J.F. STANTON, Department of Chemistry, The University of Texas at Austin, TX; M.E. HARDING and J. GAUSS, Institut für Physikalische Chemie, Universität Mainz, Germany.

The hydrogen trioxy radical is thought to be a possible sink for hydroxyl radicals in the atmosphere. The calculated structure and properties of this interesting molecule will be discussed and compared to experiment.